OpenStudy (anonymous):

Chapter 5:Indices and Logarithms @ganeshie8

OpenStudy (anonymous):

solve the following equation

OpenStudy (anonymous):

\[3\log _{2}5+\log _{2}(2x-1)-\log _{2}(x+2)=1\]

ganeshie8 (ganeshie8):

Notice \(3\log_2 (5) = \log_2 (5^3)\)

ganeshie8 (ganeshie8):

write the left hand stuff as a single log

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[\log _{2}5^3+\log _{2}2+\log _{2}x+\log _{2}-1-\log _{2}x+\log _{2}2=1\]

ganeshie8 (ganeshie8):

Hey no

ganeshie8 (ganeshie8):

below is false : \[\log_b (x+y) \ne \log_b(x) + \log_b(y)\]

OpenStudy (anonymous):

Then factorize?

OpenStudy (anonymous):

\[\log _{2}(5^3+2x-1-x-2)=1\]

ganeshie8 (ganeshie8):

\[3\log _{2}5+\log _{2}(2x-1)-\log _{2}(x+2)=1\] \[\log _{2}5^3+\log _{2}(2x-1)-\log _{2}(x+2)=1\] combine the first two logs using product property: \[\log _{2} \left(5^3(2x-1)\right)-\log _{2}(x+2)=1\]

ganeshie8 (ganeshie8):

combine the logs again using quotient property of logarithm : \[\log_2\left(\dfrac{5^3(2x-1)}{x+2}\right)=1\]

ganeshie8 (ganeshie8):

right ?

OpenStudy (anonymous):

\[\frac{ 5^3(2x-1) }{ x+2 }=2\]

ganeshie8 (ganeshie8):

yes!

ganeshie8 (ganeshie8):

cross multiply and solve x

OpenStudy (anonymous):

\[5^3(2x-1)=2(x+2)\]

OpenStudy (anonymous):

\[125(2x-1)=2x+4\]

OpenStudy (anonymous):

\[250x-125=2x+4\]

OpenStudy (anonymous):

\[248x=125+4\]

OpenStudy (anonymous):

\[x=0.5202\]

OpenStudy (anonymous):

Thnx @ganeshie8 :)

ganeshie8 (ganeshie8):

good job !